OpenStudy (anonymous):

Latanya made a 17-minute phone call from her hotel in France and was charged 17 euro. At the time, $1 was worth 0.82 euro. Find the cost per minute of the call in dollars.

OpenStudy (mathstudent55):

17 euro/17 minutes = 1 euro/minute 1 euro/minute * $1/0.82 euro

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

17? @mathstudent55

OpenStudy (anonymous):

1/.82 = x/17 ($1 is to .82 euros = $x is to 17 euros) cross multiply (.82)(x) = (1)(17) .82x = 17 x = 17/.82 x = 20.73 now that is the price for 17 minutes. You want the price for a minute. 17/20.73 = 1/x (17 min to 20.73 = 1 min to $x) cross multiply 17x = 20.73 x = 20.73/17 x = 1.219 = 1.22 It would cost $ 1.22 per minute